# ScrollTimes: Tracing the Provenance of Paintings as a Window into History

### v1

@@ 4.1.1 Inscription Processing
The inscriptions used in this work come from the HD database. We directly employ the Named Entity Recognition (NER) technique to process inscriptions’ textual contents (Fig. 3A). Specifically, NER aims to identify entities with specific meanings in the text, including person names, place names, organization names, and proper nouns. We adopt a Chinese NER model called RaNER [63], which incorporates a Transformer-CRF model with StructBERT [62] as the pre-trained model. This model can identify entities such as addresses, people, times, organizations, and other proper nouns in the inscriptions.
Nevertheless, NER models are subject to limitations regarding the length of the input text, hindering the processing of some inscription texts. To alleviate the obstacle, we deliver a sliding window approach to extract sentences of reasonable length, which are then fed into the model for entity recognition. Subsequently, a straightforward voting strategy is applied to manage the overlapping short sentence parts, culminating in the final NER classification results. To obtain additional details regarding entities like figures, times, and locations, the NER results are post-processed to associate them with open-source databases. However, the model tends to recognize boundaries that are excessively large, leading to erroneous results during the testing phase. To address this issue, a sliding window is implemented that is based on individual characters to extract sub-words, which are subsequently queried separately. If a sub-word produces a relevant result, the range of the recognition results is rectified and associated with the corresponding information. This approach effectively boosts the accuracy of entity recognition, as demonstrated by test results.
Figure 3: After being processed by various models and validated and supplemented through multiple databases, inscription (A), core painting (B), and seal (C) are organized into a series of historical events (D) for the purpose of constructing the painting biography through the ScrollTimes (E).
@@ 4.1.2 Core Painting Processing
The relationship between students and teachers among painters plays a crucial role in shaping their painting techniques and styles. To provide an insightful exploration of a large number of paintings, it is necessary to recommend similar paintings. As a handscroll usually contains a substantial amount of textual content, it is essential to extract the core painting from the handscroll before calculating similarities, ensuring a more precise and rigorous calculation of painting similarity (Fig. 3B).
In this paper, we utilize a deep neural network known as DeepLab V3 [14] to extract core painting. DeepLab V3 has achieved impressive performance in semantic segmentation within the field of natural images. It uses the ResNet [35] as the backbone for feature extraction, complemented by an Atrous Spatial Pyramid Pooling module. This module analyzes high-level features extracted from the image and generates multi-scale context features, thereby enhancing the ability to segment objects accurately. After training DeepLab V3, we employ it to extract core paintings in HD, building a core painting gallery.
To measure the similarity between core paintings, we use a self-supervised learning method called BYOL [32]. This method involves projecting images into a feature space using an encoder and computing their similarity in that space. During training, two augmented views of each image are fed into an online encoder and a momentum encoder to construct consistency constraints and optimize the encoder. We then retrieve the most similar core paintings from the gallery by comparing the cosine similarities of their query painting features.
@@ 4.1.3 Seal Processing
To detect and match the seals in the handscroll, and understand the owner and content of the seals, we employ computer vision methods for relevant processing (Fig. 3C). We employ a neural network named YOLOX [31] to detect seals in the handscroll. YOLOX integrates the benefits of a series of YOLO models, offering exceptional detection speed and accuracy. In the experiments, we curate a dataset of Dynasty Song paintings and perform annotations on the seals within the paintings. The trained model facilitates the detection of seals from handscrolls and their subsequent isolation for seal matching. The handscrolls contain numerous seals, and their attribution to specific authors or content may prove challenging. Thus, matching with available open-source seal databases becomes necessary. We collect a standard seal gallery including 36302 seal images, along with the corresponding author and textual content. By leveraging this gallery, we match the detected seals from handscrolls to issue the author and content information to the seals.
@@
We train another BYOL model with a ResNet-101 backbone for seal feature extraction. Notice that seal matching performance may be impacted by the intricate and diverse backgrounds of the handscrolls. Moreover, the gallery samples are limited, with each seal in the gallery having only one image. Therefore, we augment the gallery by gathering a set of seal backgrounds from handscrolls and randomly substituting the seal backgrounds to create a new seal dataset with several times more samples. To replace seal backgrounds, we apply Otsu’s method [51] for binary thresholding, which is efficient and effective. With the trained BYOL, we project seal images into feature vectors for seal matching. For each detected seal from handscrolls, we employ a vector retrieval algorithm, specifically Local Sensitive Hashing [38], to search for the most similar seals from the gallery, thereby accelerating the retrieval process.
4.2 Data Fusion
In Section 4.1, we extracted entities, including time, location, figure, and thing from inscriptions core painting, and seals, but the accuracy and consistency of the information have not been validated. To ensure a clear and well-structured account of events, we cross-check and enrich the data by referring to multiple databases (Fig. 3D) (T3, T4).
Figure validation supplement. To identify the figure entities, we query PerAD and CBDB to validate and supplement the information. However, the identification of the individuals can be challenging due to the extensive use of aliases and personal names with prefixes in ancient Chinese. To address this issue, we adopt the following approaches:
@@ 4.2 Data Fusion
  • Directly querying PerAD and CBDB. Although CBDB has comprehensive descriptions of historical figures, PerAD provides valuable data, especially related to religious figures and deities. For example, the painter Zhao Mengfu can be queried in both databases, but the named entity Jin Mu can only be found in PerAD as an alternative name for Queen Mother of the West [10].
  • Querying with sliding-window segments. For instance, “Han Lin Qian Pu”(“Hanlin” was an academic and administrative institution of higher learning founded in China) can be segmented and corrected as Qian Pu to supplement the information of the corresponding figure.
  • Priority determination for figures with the same name based on the era and social relationships. For example, an inscription from the Yuan Dynasty mentions a figure named Gong Jin Fu, where “Fu” is an honorific for a man. Despite there being more than twenty historical figures with the name Gong Jin, only Zhou Mi and Li Kezhong is from the Yuan Dynasty, with 47 and 3 social relationships, respectively. Therefore, we can conclude that the figure referred to is Zhou Mi.
  • Allowing art historians to interactively choose the correct figure. While automated methods are helpful, they may not guarantee accuracy in identifying historical figures. Therefore, art historians can conduct fuzzy searches on figure names and select the correct figure based on the presented list and their prior knowledge.
@@
Location validation. Place entities have fewer errors than figure names, yet challenges still emerge due to the prevalence of aliases, prefixes, and suffixes. Analogous to the procedure for figures, we cross-reference PlaAD and CBDB databases to rectify place entities. For instance, the toponym Qizhou can be recognized by both databases, while Que mountain can only be found in PlaAD.
@@
Time standardization. The model’s recognition of time entities is generally precise, with the primary task of converting them into Gregorian calendar years. In ancient China, years are typically indicated by the era name followed by either the year number under that era or the Sexagenary cycle [55], a traditional time reckoning system. Regular expression matching is used to extract the corresponding era name, numeric value, and Sexagenary cycle name, and the era’s start and end dates are queried to calculate the exact year. For example, if the start year of the “Yuanzhen” era is 1295, then the time entity Yuanzhen second year corresponds to the year 1296. Qianlong Wuchen denotes the year 1748, as the “Qianlong” era ranges from 1736 to 1796, with 1748 being the “Wuchen” (a Sexagenary cycle name) year.
Thing supplement. Due to the diverse entity types, such as mountains, birds, and paintings, automatic methods face difficulties verifying and supplementing them. Art historians’ prior knowledge is required to supplement events with entities. For example, if a CBDB records that someone wrote inscriptions in paintings, and the art historian adds the entity “Autumn Colors on the Que and Hua Mountains” to the event, it constitutes a complete event: someone inscribes an inscription on“Autumn Colors on the Que and Hua Mountains” painting. As a quantitative analysis, we examine the performance of NER on a separate test set. The post-processed NER results achieve an F1 score of 0.558, significantly surpassing the F1 score of 0.483 yielded by the vanilla NER results.
@@ 5 Visual Analytic System
Figure 4: The original painting (A) is transformed into a circular ring shape (B). The seam carving algorithm is used to remove white space (C). After preserving the core painting, reducing silk, and repairing text distortion, the final representation is displayed (D). * To better visualize the silk texture, we have applied visual enhancement to the silk in this caption.
We propose the visual analysis system ScrollTimes to display the features of the handscroll and support art historians in exploring interested figures interactively, helping art historians construct painting biography.
5.1 Handscroll Features View
The Handscroll Features View offers art historians an accessible way to browse handscrolls and gain insights from feature visualization (Fig. 3E1) (T1, T2). The redesign of the handscroll considered two principles: 1) conserving screen space to emphasize the crucial points of the scroll (Section 5.1.1); 2) conforming to contemporary reading habits by modernizing the ancient text reading form (Section 5.1.2).
5.1.1 Handscroll Circle Design
Transforming the handscroll into a circle. To maximize the efficient use of system space for handscrolls with wide aspect ratio disparities (Fig. 4A), we transform them into a circular ring shape (Fig. 4B).
@@ 5.1.1 Handscroll Circle Design
The handscroll is divided into two equal-width parts and projected into a 180-degree semicircular ring, which is then flipped and concatenated to form a complete circular ring. To achieve the desired circular effect, each part of the original image is divided into rectangular blocks and proportionally projected into trapezoids based on pixel positions.
Processing white space. We use seam carving [9, 22] algorithm for handscroll white space processing (Fig. 4C). This involves scaling the original handscroll while preserving high-energy content and removing low-efficiency information. Seam carving constructs an energy map of the painting by calculating the discrete pixel gradients. A seam is defined as a connected path of pixels running from the top to the bottom of the painting. Using dynamic programming, the algorithm finds and removes the connected path with the lowest cumulative energy.
@@
Regulating the ratio of the core painting. In the handscroll, the core painting contains more content and has higher information density than other parts. During interviews with art historians, it is determined that the core painting should occupy one-third of the entire length of the handscroll. We resize the handscroll to a fixed length, with one-third being the extracted core painting (Fig. 4D) and the remaining filled with other parts. In addition, some texts are severely deformed due to excessive compression. To solve this problem, we divide non-core painting parts into smaller blocks and reduce the compression ratio for blocks with over-deformable texts (Fig. 4D).
@@
Interaction. We provide three interactive ways for art historians to observe the parts of their interests in (Fig. 4D). 1) [Uncaptioned image] When the mouse moves over the handscroll, the original image of the current section is displayed. 2) [Uncaptioned image] When the mouse slides over the blank space of the circular ring, the ring can be rotated accordingly. 3) [Uncaptioned image] The original handscroll is provided below the circular ring, which supports the art historians in selecting a part of the area. The perspective center of the circular ring will rotate to the selected area. Art historians also can zoom in on the original painting to view the details of the handscroll [Uncaptioned image] .
5.1.2 Handscroll Features Word Cloud
Seals and inscriptions on handscrolls are not systematically distributed and often pose reading challenges for art historians due to the ancient Chinese writing style. To make them more accessible, we apply a word cloud visualization technique to convert ancient writing into a modern format, which facilitates easy access to these features while aligning with modern reading habits. In Section 4.1.3, we extracted the sealers on the handscroll and simplified their visual form into squares (Fig. 3E1-1). The text in the middle of the square represents the sealer’s surname, while the transparency of the red background of the text encodes the number of seals they left on this scroll. This allows art historians to quickly understand which sealer left the most seals on the painting scroll. The color shade of the outer frame reflects the number of seals they left on the entire handscroll dataset, providing insight into their importance in the history of collecting and appreciation (Fig. 5). To reflect the changes in dynasty during the transmission of the painting, we cluster the sealers according to their respective dynasties (Fig. 3E1-1). In Section 4.1.1, we conducted entity extraction on the inscriptions and generated word cloud visualizations based on entity classification ([Uncaptioned image]Time, [Uncaptioned image]Location, [Uncaptioned image]Figure, [Uncaptioned image]Thing). We use font color depth to encode the frequency of the words (Fig. 3E1-2). Upon clicking on a particular sealer, the solid lines on the view illustrate the distribution of their seals on the handscroll, while the dotted lines indicate the entities mentioned in their inscriptions. Additionally, the Seals and Inscriptions View (Fig. 3E1-3) provides a comparison between the seals on the handscroll and the images in the seal database, along with the original inscriptions that have been identified and labeled by the computer based on entity classification. This view is designed to assist art historians in identifying and revising information based on prior knowledge (T3).
5.2 Linked Data View
To provide supplementary information for the handscroll and enable art historians to have a more comprehensive understanding of the painting and conduct relational analysis, we visualize the linked data from other databases (Fig. 3E2) (T4). We select the ego network relationship diagram, which is familiar to art historians, to visualize figure social relationships (Fig. 3E2-1). Additionally, we also present the relationships between cohorts through matrix visualization (Fig. 3E2-2).
5.2.1 Ego Network View
The Ego Network View provides a social network analysis that focuses on a single historical figure. If an art historian is interested in a particular figure from a seal or inscription (Fig. 3E2-1), they can obtain additional information through this view. We encode the data related to the paintings using a hairbrush icon ([Uncaptioned image]Painter) to represent the painter, with the size of the icon indicating the number of paintings associated with that painter. The figure’s basic information (e.g. name, birth year, death year, nickname, status) will be displayed when the art historian hovers over the figure icon, while all of their events will be shown in the table below. For the types of historical events that facilitate further analysis by art historians, we have coded them separately with different colors (Fig. 5). Art historians can click on the lines connecting the figures to view the events associated with them in the table below. The thickness of the lines represents the number of events, providing a visual indication of the strength of the relationship between the figures.
5.2.2 Cohort View

### v2

@@ 4.2.1 Inscription Processing
The inscriptions used in this study are obtained from the handscroll database. Our objective is to extract valuable information from these inscriptions, including Time, Location, Thing (various objects, such as mountains, birds, and paintings), and Figure (T1, T3) (Fig. 3A1).
We evaluate four Chinese Named Entity Recognition (NER) pre-training models, namely BERT-WWM, GuwenBERT, SikuBERT, and RaNER [75], to identify entities. We select the RaNER model due to its excellent performance in extracting entities from our dataset. However, the RaNER model has limitations when it comes to processing lengthy input texts. To overcome this limitation, we implement a sliding window approach [48] to extract sentences of a manageable length. These sentences are then input into the model for entity recognition. Following this, we employ a simple voting strategy to handle overlapping segments of shorter sentences, ultimately producing the final RaNER classification results.
Figure 3: Handscroll elements, including (A1) inscriptions, (A2) core paintings, and (A3) seals, undergo processing by various automated models to extract entities. Subsequently, (B1) these entities are validated and supplemented through multiple databases. Finally, (B2) we compile historical figures and events in chronological order.
@@ 4.2.2 Core Painting Processing
The extraction of the core painting is a crucial aspect of understanding the visual properties of a handscroll (T1). Furthermore, since painting places great emphasis on copying and transferring styles [15], the inheritance of painting styles can be traced through comparisons of similarities, akin to viewing the “genealogy” of a painting (i.e., other painters that were influenced by it) (T3) (Fig. 3A2).
In this paper, we employ a deep neural network known as DeepLab V3 [13] due to its remarkable performance in semantic segmentation within the field of natural images for extracting the core painting. This method utilizes ResNet [39] as the backbone for feature extraction. To enhance the accuracy of object segmentation, the approach analyzes high-level features extracted from the image and generates multi-scale context features using an atrous spatial pyramid pooling (ASPP) [37] module. Following the training of DeepLab V3, we utilize it to extract core paintings in the HD dataset, creating a core painting gallery.
To measure the similarity between core paintings, we use a self-supervised learning method called BYOL [33]. This method involves projecting images into a feature space using an encoder and computing their similarity in that space. During training, two augmented views of each image are fed into an online encoder and a momentum encoder to construct consistency constraints and optimize the encoder. We then retrieve the most similar core paintings from the gallery by comparing the cosine similarities of their query painting features. Then, we obtain core paintings and related painters.
@@ 4.2.3 Seal Processing
Seals play a pivotal role in tracing the provenance of paintings. Our approach to handling seals encompasses seal detection, extraction, and matching. Ultimately, we construct a seal dataset that includes seal images and sealers (T1, T3) (Fig. 3A3).
@@
For seal detection, we employ a neural network named YOLOX [32] to detect seals in the handscroll. YOLOX integrates the benefits of a series of YOLO models, offering exceptional detection speed and accuracy. In the experiments, we curate a dataset of Song Dynasty paintings and perform annotations on the seals within the paintings. The trained model facilitates the detection of seals from handscrolls and their subsequent isolation for seal matching.
@@
For seal matching, we collect a standard seal gallery including 36302 seal images, along with the corresponding author and textual content [2]. By leveraging this gallery, we match the detected seals from handscrolls to identify the sealers and content information. We train another BYOL model with a ResNet-101 backbone for seal feature extraction. Notice that seal matching performance may be impacted by the intricate and diverse backgrounds of the handscrolls. Moreover, the gallery samples are limited, with each seal in the gallery having only one image. Therefore, e augment the gallery by gathering a set of seal backgrounds from handscrolls and randomly substituting the seal backgrounds with for binary thresholding [62]. After finetuning the BYOL with the new dataset, we project seal images into feature vectors for seal matching. We employ Local Sensitive Hashing (LSH) [42] to search for the most similar seals from the gallery efficiently.
4.3 Context Information Extraction
At the contextual level, it is necessary to build a social network that involving historical figures and related historical events (R2). We extracted entities from inscriptions, core painting, and seals (Section 4.2). In order to establish a robust social network, it is essential to validate these entities (T3) through the use of multiple databases. Additionally, to enrich contextual information, we need to integrate data from multiple databases (T4) (Fig. 3D).
Figure Validation and Supplementation. To validate and enhance the identification of historical figures, we cross-reference the PerAD and CBDB databases and show the data provenance as explanations of automated methods [29]. Nevertheless, owing to the prevalence of aliases and the use of prefixed names [81], such as official titles or ancient Chinese locations, certain entity recognition results cannot be directly mapped to the names in the database. To tackle this challenge, we employ the following strategies:
@@ 4.3 Context Information Extraction
  • Querying PerAD and CBDB directly. While CBDB provides extensive descriptions of historical figures, PerAD contains valuable data related to religious figures and deities. For example, the painter Zhao Mengfu can be found in both databases, but the named entity Jin Mu is exclusively available in PerAD, where it functions as an alternative name for the Queen Mother of the West [10].
  • Querying with sliding-window segments. For instance, “Han Lin Qian Pu” cannot be directly queried. Therefore, by sliding the window in reverse order, it was split into the following words: “Lin qian Fu,” “Han lin Qian,” “Qian Fu,” “Lin qian,” “Han lin.” After searching in order, it was found that “Qian Fu” can be queried, so the record matching result is Qian Pu.
  • Determining priority for same name. When encountering duplicate names in the query results, we employ a selection process guided by art historians. This process involves two primary rules: First, Era-Based selection. For instance, an inscription from the Yuan Dynasty mentions a figure named Gong Jin, and there are more than twenty historical figures in the database with the alias Gong Jin, we prioritize figures who lived during the Yuan Dynasty, such as Zhou Mi and Li Kezhong. Second, Social Relationship criteria. For instance, Zhou Mi has 23 social connections with other confirmed figures mentioned in the painting, whereas Li Kezhong has no such relationships. We identify the figure as Zhou Mi due to the stronger network of associations.
  • Allowing interactive figure selection. While automated methods are beneficial, they may not ensure precise historical figure identification. Hence, art historians can perform fuzzy searches on names and make selections based on the provided list and their expertise.
@@
Location Validation. Challenges arise due to aliases, prefixes, and suffixes in place names. Similar to the historical figures’ process, we cross-reference PlaAD and CBDB databases for matching place entities. For instance, Qizhou is found in both databases, while Que mountain is only available in PlaAD.
@@
Time Standardization. In ancient China, time are typically indicated by the era name followed by either the year number under that era or the Sexagenary cycle [65], a traditional time reckoning system. Regular expression matching is used to extract the corresponding era name, numeric value, and Sexagenary cycle name, and the era’s start and end dates are queried to calculate the exact year. For example, if the start year of the “Yuanzhen” era is 1295, then the time entity Yuanzhen second year corresponds to the year 1296. Qianlong Wuchen denotes the year 1748, as the “Qianlong” era ranges from 1736 to 1796, with 1748 being the “Wuchen” (a Sexagenary cycle name) year.
Thing Supplement. Due to the various entity types, such as mountains, birds, and paintings, automated methods encounter challenges in verifying and complementing them. Art historians’ expertise is essential for augmenting events with these entities. For instance, if a record in CBDB mentions that someone wrote inscriptions for the painting “Autumn Colors on the Que and Hua Mountains,” and an art historian includes the entity in the event, it results in a comprehensive event: Qian Pu inscribed an inscription on the painting “Autumn Colors on the Que and Hua Mountains” in 1446.
NER Performance Analysis. To assess the impact of post-processing through database matching on result accuracy, we conducted experiments by comparing the entities obtained after database matching as corrected entities with those identified by the NER process. We annotated a subset of the dataset for testing and evaluated the results using the F1 score as a performance metric. The post-processed NER results achieved an F1 score of 0.558, which significantly outperformed the F1 score of 0.483 obtained from the basic NER results. This experiment demonstrates that post-processing through database matching effectively enhances result accuracy.
@@ 5 Visual Analytic System
Figure 4: (A) The original painting is transformed into (B) a circular ring shape. (C) The seam carving algorithm is used to remove white space. After preserving the core painting, reducing silk, and repairing text distortion, (D) the final representation is displayed. * To better visualize the silk texture, we have applied visual enhancement to the silk in this caption.
In this section, we propose the visual analysis system, ScrollTimes, which comprises three-tiered components to facilitate the analytical tasks outlined in Section 3.2.
5.1 Artifact-Level Component
The Artifact-Level Component (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA) visualizes the features (T1) and correlations (T2) of handscroll elements, making it accessible for art historians to browse handscroll and gain insights (R1).
5.1.1 Summarizing Handscroll Element Features
Seals and inscriptions on the handscroll are typically not systematically distributed, often posing reading challenges for art historians due to the ancient Chinese writing style. To enhance accessibility, we employ a word cloud visualization technique (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA1, A2) to modernize ancient writing, facilitating user access while accommodating contemporary reading practices (T1). Furthermore, ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA3 provides comprehensive details about seals and inscriptions.
The Seal Word Cloud View. In Section 4.2.3, we extracted the sealers on the handscroll and simplified their visual form into squares (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA1, Fig. 5A1). The text in the middle of the square represents the sealer’s surname, while the transparency of the red background of the text encodes the number of seals they left on this handscroll. This allows art historians to quickly understand which sealer left the most seals on the painting scroll. The color shade of the outer frame reflects the number of seals they left on the entire handscroll dataset, providing insight into their importance in the history of collecting and appreciation. To reflect the changes in dynasty during the transmission of the painting, we cluster the sealers into word cloud visualizaton according to their respective dynasties (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA1).
The Inscription Word Cloud View. In Section 4.2.1, we performed entity extraction on the inscriptions. We utilize word cloud visualizations that are systematically categorized by various types of entities, including [Uncaptioned image]Time, [Uncaptioned image]Location, [Uncaptioned image]Figure, and [Uncaptioned image]Thing. The frequency of words is encoded and represented through discernible variations in font color depth (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA2).
The Seal and Inscription Detail View. This view (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA3) provides a comparison between the seals on the handscroll and the images in the seal database, along with the original inscriptions that have been identified and labeled by the computer based on entity classification.
5.1.2 Demonstrating Handscroll Element Correlations
Art historians require an accessible method for exploring handscroll and gaining insights from correlations among elements (T2). We have redesigned the handscroll in a circular format and provide an interactive means of associating elements.
The Handscroll Circle View. This view (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA4) redesign handscroll considering two principles: 1) displaying the relationships among elements; 2) conserving screen space to highlight the key aspects of the handscroll. Specifically, We demonstrate the design process as follows:
@@ 5.1.2 Demonstrating Handscroll Element Correlations
  • Transforming the handscroll into a circle. The handscroll has wide aspect ratio disparities (Fig. 4A) that hinder fast retrieval of related contents. To maximize the efficient use of system space, we transform them into a circular ring shape (Fig. 4B). The handscroll is divided into two equal-width parts and projected into a 180-degree semicircular ring, which is then flipped and concatenated to form a complete circular ring. To achieve the desired circular effect, each part of the original image is divided into rectangular blocks and proportionally projected into trapezoids based on pixel positions.
@@
  • Regulating the ratio of the core painting. In the handscroll, the core painting contains more content and has higher information density than other parts. During interviews with art historians, it is determined that the core painting should occupy one-third of the entire length of the handscroll. We resize the handscroll to a fixed length, with one-third being the extracted core painting (Fig. 4D) and the remaining filled with other parts. We adapt the seam carving [9, 21] algorithm for handscroll white space processing (Fig. 4C). This involves scaling the original handscroll while preserving high-energy content (i.e., texts and drawings) in the handscroll and removes low-efficiency information (i.e., white space). However, seam carving heavily removes the spaces between text characters and distort the appearance (Fig. 4C). To solve this problem, we divide non-core painting parts into smaller blocks and reduce the compression ratio for blocks with over-deformable texts (Fig. 4D).
@@
  • Interaction. To aid art historians in focusing on areas of interest from a central perspective, we offer two interactions. First, [Uncaptioned image] the original handscroll is displayed below the circle view, allowing art historians to select a specific area. The central perspective of the circle view will adjust accordingly (Fig. 4A). Second, [Uncaptioned image] when the mouse hovers over the blank space of the circle view, the ring can be rotated accordingly (Fig. 4D). Additionally, to connect the painting and its elements, we provide interactive lines. Upon clicking on a particular sealer, the solid lines on the view illustrate the distribution of their seals on the handscroll, while the dotted lines indicate the entities mentioned in their inscriptions (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA5).
Justification. Beyond its historical significance, the handscroll is a cherished art form. Art historians, in their reverence for these traditional works, question the circular transformation, which does not align with their mental models and raise concerns about potential damage. However, when we introduce our handscroll circle design, which efficiently analyzes the characteristics of the handscroll, particularly in connecting scattered elements on the handscroll, art historians greatly appreciate its advantages. Additionally, to cater to art historians’ appreciation for the original artwork, we offer two interactive options: hovering over the handscroll[Uncaptioned image] displays the original image of the current section (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA6), and clicking the zoom-in button [Uncaptioned image] to view details (Fig. 4A).
5.2 Contextual-Level Component
To provide supplementary information for the handscroll and enable art historians to have a more comprehensive understanding of the painting and conduct relational analysis (R2), we extract historical figures and events (Section 4.3) and present the data through a linked data view (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB) (T3, T4).
5.2.1 Exploring Ego Social Networks
We use the ego network relationship diagram (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB1), a familiar diagram for art historians, to display the social relationships of individual historical figures (T3, T4). To help art historian quickly access painting information, we use a hairbrush icon ([Uncaptioned image]Painter) to represent painters, with the icon’s size corresponding to the number of associated paintings. User can click on a specific historical figure mentioned in a seal or inscription (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA1-A3) to access basic information. Hovering over a figure icon displays their name, birth and death years, nickname, and status, along with associated events in a table below. Clicking on connecting lines between historical figures reveals related events in the table, with line thickness indicating relationship strength.
5.2.2 Exploring Cohort Social Networks